# Impact of the Hole Gas on Optically Detected Magnetic Resonance in (Cd,Mn)Te Based Quantum Wells

### v1

@@ Abstract
Optically detected magnetic resonance (ODMR) is a useful technique for studying interactions between local spins (magnetic ions) and carrier gas. We present the ODMR study of single (Cd,Mn)Te/(Cd,Mg)Te quantum wells (QWs) with the hole gas. We observe different characteristics of the ODMR signals obtained simultaneously using the optical signals of neutral and positively charged exciton. From that, we infer an existence of local fluctuations of carrier gas density resulting in separate populations of Mn2+  ions. At the same time, the shape of the ODMR signal contains information about the temperature of the magnetic ions involved in the absorption of the MW. Studying it in detail provides even more information about the interactions with charge carriers. In the QW, two separate ensembles of ions are thermalized differently in the presence of carriers.
@@ I INTRODUCTION
In the QWs being the subject of the research presented here, it is possible to control the concentration of carriers by using the additional illumination Kossacki et al. (1999); Maślana et al. (2003); Kossacki (2003); Kossacki et al. (2004); Aku-Leh et al. (2007). Using the illumination above the energy of the barrier results in increase of holes concentration. While the sample is illuminated only below the barrier energy – the quantum well is more neutral. This effect can be observed in both photoluminescence and reflection measurements. This additional above-bandgap illumination is usually realized with the excitation laser for the photoluminescence study. During reflectivity measurements, apart from incident light in the range corresponding to exciton structures, the sample should be illuminated additionally above the barrier energy gap – to increase carrier density in the QW. Moreover, carrier concentration can be estimated based on the ratio of the amplitudes of the spectral lines corresponding to the neutral X   and the charged X+   exciton present in the optical spectrum measured in a zero magnetic field. That allows one to simultaneously change and probe the carrier concentration in the nanostructure on an optical basis.
@@
II Samples and characterization
The studied structures were single 10 nm (Cd, Mn)Te QWs grown by molecular beam epitaxy (MBE) between symmetrical 50-nm (Cd,Mg)Te barrier layers with Mg content of about 20 %. They were prepared on GaAs substrate with a thick (5500 nm to 6000 nm) CdTe buffer layer. The Mg content in the barrier layer was determined by the reflectance measurements from the bandgap energy of the barrier Waag et al. (1993). The Mn2+ content in the QW material close to 0.3 % was chosen to ensure sufficiently large Zeeman splitting while preserving narrow excitonic lines. The exact value was confirmed by fitting the Brillouin function Gaj et al. (1994) to the exciton Zeeman shift obtained from reflectance measurements in a magnetic field, see Fig.1(b).
Generally, (Cd,Mn)Te/(Cd,Mg)Te QWs are naturally p-type, even without intentional doping. In such a case, hole gas originates from background doping of the barrier and/or from the surface states, which can be especially influential for the QW placed near the surface Maślana et al. (2003). Indeed, the presented samples contain hole gas in the QW. The sign of the carriers can be confirmed by using measurements of the photoluminescence signal in the magnetic field and observation of the spin singlet-triplet transition Kossacki et al. (2004); Łopion et al. (2020). The same group of samples was presented in Łopion et al. (2020), where their optical properties (along with the field of the spin singlet-triplet transition) were studied as a function of external illumination.
@@ II Samples and characterization
In the magnetic field, both exciton lines show a giant Zeeman shift. One can consider how introducing carriers into the system affects the thermalization of magnetic ions excited by microwave radiation in the ODMR experiment. Figure 1(b) shows the measurement of giant Zeeman splitting as a function of the magnetic field with low carrier density (without additional illumination). The modified Brillouin function was fitted to the splitting of exciton lines X   in a magnetic field. From that fit, the temperature of the system of polarizing ions is estimated at 1.73 K. An analogous measurement was carried out for the increased carrier density – the effective temperature of the ions increased slightly in this situation, up to 1.81 K.
Figure 1(c) shows reflection spectra measured with increased carrier density for different magnetic fields. In one circular polarization, the X+   line disappears. The positively charged exciton consists of two holes with opposite spins. The optical transition X+   associated with the formation of a hole with a specific spin depends on the prior existence of a hole with an opposite spin Huard et al. (2000). Hence, the oscillator strength of the charged exciton X+   is proportional to the concentration of holes in the opposite spin subband. By tracking the changes in the relative intensity of the reflection lines X   and X+, it can be seen that in this case, the hole gas is completely polarized already in the field of about 0.1 T, Fig. 1(d). The following function, derived from Boltzmann distribution of holes between two spin levels, was fitted to the intensity changes in the field presented in the graph:
@@
where $A_{\mathrm{max}}$ is the maximum intensity of a given line, $\Delta_{v,Z}$ giant Zeeman splitting of the valence band, $T_{\mathrm{h}}$ is the carriers’ temperature. This results in a temperature of the carriers slightly higher than the temperature of the ion system without microwave radiation. The magnetic fields used for ODMR measurements are larger for all the cases presented in this paper. Thus, it can be assumed that the ions interact with the polarized hole gas in our experiments.
Figure 1: (a) Difference in the reflection spectrum in zero magnetic fields, in the absence and the presence of additional illumination with a blue LED. (b) Zeeman splitting of the neutral exciton line for the unilluminated situation (black points) with the fitted modified Brillouin function (gray curve) and the sample subjected to additional illumination (increased carrier density) - yellow points and yellow curve, respectively. The temperature of the ion system was a fitting parameter and was set to 1.73 K in the non-illuminated situation and 1.81 K in the illuminated situation. The black curve represents the ODMR signal obtained without the blue LED backlight for the frequency of microwave radiation $f$ = 12.9 GHz, the red curve symbolizes the modified Brillouin function for the remaining parameters exactly as in the case of the gray curve but the temperature set to 2.68 K. (c) The dependence of the reflection spectrum for different magnetic fields with increased carrier density; (d) Relative intensity of the lines of the neutral X   and positively charged exciton X+   depending on the magnetic field with fitted relationships – according to the formula (1).
@@ III ODMR - setup and method of the measurements
The transitions between Mn2+ energy levels due to absorption of the applied microwave (MW) radiation lead to a decrease in magnetization. This can be evidenced by the decrease of the giant Zeeman shift. The effect can also be expressed in terms of increased effective spin temperature, which will be discussed later in this work. All the ODMR measurements presented in this work were conducted in the pulsed mode of MW radiation and optical excitation to minimize non-resonant thermal drifts. Here, we define the ODMR signal as the difference in the energy position of the excitonic line in the reflectance spectrum for microwave and light pulses overlapping (ON) and shifted (OFF). The duration of the MW pulse is about a few milliseconds. A representative ODMR signal as a function of magnetic field (ODMR spectrum) for fixed microwave frequency is present in Fig. 1(b). Careful study of ODMR spectra can provide information about strain, what was discussed in detail by Bogucki et al. (2022). Moreover, the ODMR spectrum can provide information about the impact of interactions with carriers, which is the core of the work presented in this paper.
The amplitude of the ODMR signal (shift of the excitonic line) corresponds directly to the temperature of the magnetic ions, which create net magnetization. After the sample is excited with pulsed microwave radiation of frequency of 12.9 GHz, a decrease in Zeeman splitting for a resonant magnetic field of 0.46 T can be observed, see Fig.1(b). This reduction in Zeeman splitting corresponds to an increase in the temperature of the ion system to 2.7 K as shown by the red curve in the figure.
An important detail is that the measurement conditions are set to avoid non-resonant signals. It can be evidenced by the fact that outside the resonance pattern, the position of the excitonic line with and without microwaves is the same – see inset of Fig. 1(b). In some cases, the offset can be seen. It is connected to the non-resonant heating of the sample, Fig. 2(a). The overall temperature increase results from the resonant and non-resonant heating of the magnetic ions system, Fig. 2(b).
@@ IV Shape and position of the ODMR signal
The ODMR spectrum has a specific shape determined directly by the structure of Mn2+ energy levels. They can be described by the spin Hamiltonian:
@@
Figure 2: (a) ODMR signal power dependence, both amplitude of the signal and the offset is increasing with the applied MW power (b) Temperature derived from Zeeman splitting change vs. MW excitation power dependence. Full Giant Zeeman splitting for this sample, bath temperature, and the magnetic field are about 3 meV, marked with a red line. The temperature of magnetic ions derived from Zeeman splitting measured without MW radiation is marked. The MW excitation power in the ODMR measurements was set to -10 dBm to avoid non-resonant signals, which are represented by increase of the offset.
@@
Figure 3: (a) Reflection spectrum of this sample measured at temperature $\sim$1.6 K in magnetic field $B=0.463$ T with microwave radiation resonant for this field (12.9 GHz) at in the "ON" and "OFF" states, the change of position is visible on both visible exciton complexes; (b) The ODMR signal as the difference in the positions of the energy lines observed in the reflection spectrum in the presence and absence of microwave radiation at the frequency $f$ = 12.9 GHz depending on the value of the applied external magnetic field; the signal observed on the line of the charged exciton CX   is shifted towards lower magnetic fields compared to the signal obtained on the neutral exciton X. o $\Delta B_{\mathrm{res}}$ = 14 mT
$H_{\mathrm{Mn^{2+}}}=g_{Mn}\mu_{\mathrm{B}}\boldsymbol{B}\cdot\widehat{\boldsymbol{S}}+A\widehat{\boldsymbol{I}}\cdot\widehat{\boldsymbol{S}}\\ +\frac{a}{6}\left(\widehat{S}_{x}^{4}+\widehat{S}_{y}^{4}+\widehat{S}_{z}^{4}-\frac{S(S+1)(3S^{2}+3S-1)}{5}\right)\\ +D\left(\widehat{S}_{z}^{2}-\frac{S(S+1)}{3}\right),$ (2)
@@
The first term of the Hamiltonian describes the Zeeman splitting of energy levels in a magnetic field due to the electrons d-shell spin ($S$). $g_{\text{Mn}}$ is the g-factor of the manganese ion. The second term describes hyperfine splitting, where $\widehat{\boldsymbol{I}}$ denotes the manganese ion nuclear spin operators. Mn2+  ion has electronic spin S = 5/2 and nuclear spin I = 5/2. The third term for the level-splitting part of the Hamiltonian due to the placement of the ion in the crystal field with the symmetry Td.
@@
The last term refers to zero field splitting related to lowering the symmetry of the crystal and corresponds to the uniaxial stress associated with deformation in the $z$ direction. The stress-induced parameter of the axial component of the crystal field $D$ is a parameter characteristic for a given sample, resulting from the parameters of growth, composition and thickness of individual layers. $D$ can be determined from the mismatch of the material lattice constants of the quantum well $a_{\mathrm{QW}}$ and the buffer $a_{\mathrm{buf}}$ by determining the strain $\varepsilon=\frac{a_{\mathrm{QW}}-a_{\mathrm{buf}}}{a_{\mathrm{QW}}}$. Finally, the $D$ parameter Qazzaz et al. (1995) can be determined: $D=-\frac{3}{2}G_{11}\left(1+\frac{2C_{12}}{C_{11}}\right)\varepsilon$, where $G_{11}$ = 72.272.272{.}272.2 neV and $\frac{2C_{12}}{C_{11}}=0.7$Bogucki et al. (2022); Greenough and Palmer (1973).
@@
The electronic and nuclear spin coupling for S and I equal to 5/2 results in 36 energetic states of Mn2+ ion. In a zero magnetic field, the splitting of those states corresponds mainly to the strain ($D$ parameter). For low magnetic fields ($B<0.05$ T), the evolution of the states is complicated, and the eigenstates of the Hamiltonian are still not eigenstates of the $S_{z}$ operator. The Zeeman term dominates Hamiltonian for higher magnetic fields, and the energy levels begin to have a well-defined spin projection to magnetic field quantization $z$ direction. Considering the selection rules for magnetic dipole transitions ($\Delta S_{z}=\pm 1$, $\Delta I_{z}=0$), the ODMR spectrum contains 6 sets with 5 transitions in each.
@@
The shape of the ODMR spectrum depends on the separation and the population of energy levels participating in microwave absorption. The intensity of a single transition corresponds to the occupation of the states between which the transition occurs—thus, it is strictly connected to the temperature of the magnetic ions. Qualitatively, the features observed in the higher magnetic field correspond to the transitions between lower energy levels than features observed at the lower magnetic field.
@@
For situations without additional illumination the carrier density is low and the neutral exciton transition dominates the reflectance spectrum. When additional illumination is set on the sample the carrier density is increased. That is evidenced in the reflectance spectrum, which shows another excitonic transition in the energy below X: charged exciton (X+). Both excitonic lines observed in the reflectance are sensitive to microwave radiation, Fig. 3(a). We can see that both X and CX lines move towards lower energies when the light and microwave pulses coincide in time ("ON" state) relative to the position of the lines when the light and microwave pulses are significantly temporarily separated ("OFF" state).
@@
Figure 3(b) shows in detail the ODMR signal measured as a shift of both X   and CX   lines. In the ODMR signal measured at the neutral exciton line, many narrow features can be seen related to the transitions between different states of the manganese ion split by the magnetic field, hyperfine interaction and strain. At the same time, we can notice that such a structure is not visible in the signal monitored on the charged exciton line and the shape of the signal seems to be smoother. Moreover, the maximum of the ODMR signal for CX   line is shifted relative to the maximum of the signal measured on the neutral exciton towards lower magnetic fields. These differences suggest that the two lines probe different populations of Mn2+  ions, exhibiting different paramagnetic resonance broadenings and shifts. The shift of the resonance observed on the charged state line results from interactions between magnetic ions and charge carriers. The $\Delta B_{\mathrm{res}}$ is connected to the density of carriers. Here, and in a similar situation in the following parts of the paper, the center of the resonance for the excitonic state is calculated as the center of mass of the ODMR spectrum (ODMR signal vs. magnetic field). That approach gives clear information about the position of the resonance when it is more complicated in shape, however it is important for the calculation to subtract the offset of the ODMR spectrum properly. A very similar result can be obtained by simply fitting a peak function (e.g., Gaussian) to the data.
@@
It should be noted here that both signals are observed when measuring the optical spectrum simultaneously. The ODMR signal obtained by observing the X   line appears to be unaffected by the interaction with the charge carriers, while that obtained by observing the CX   line is apparently modified. This result suggests that the absorption associated with the neutral exciton occurs in different, spatially separated regions than the absorption associated with the charged exciton. At the same time the spot of light used for reflectance measurements covers and probes both types of areas of different local carrier density. In the scale above the laser spot size ($\sim$100 $\mu m$), the sample is rather homogenous in terms of the optical response. That suggests the discussed inhomogeneities of the carrier density distribution are relatively small, below the spatial resolution of this optical measurement.

### v2

@@ Abstract
Optically detected magnetic resonance (ODMR) is a useful technique for studying interactions between local spins (magnetic ions) and carrier gas. We present the ODMR study of single (Cd,Mn)Te/(Cd,Mg)Te quantum wells (QWs) with the hole gas. We observe different characteristics of the ODMR signals obtained simultaneously using the optical signals of neutral and positively charged exciton. From that, we infer the existence of local fluctuations of carrier gas density resulting in separate populations of Mn2+ ions. At the same time, the shape of the ODMR signal contains information about the temperature of the magnetic ions involved in the absorption of the microwaves. Studying it in detail provides even more information about the interactions with charge carriers. In the QW, two separate ensembles of ions are thermalized differently in the presence of carriers.
@@ I INTRODUCTION
In the QW being the subject of the research presented here, it is possible to control the concentration of carriers by using the additional illumination Kossacki et al. (1999); Maślana et al. (2003); Kossacki (2003); Kossacki et al. (2004); Aku-Leh et al. (2007). Using the illumination above the energy of the bandgap of the barrier results in increase of the hole concentration. While the sample is illuminated only below the barrier bandgap energy – the quantum well is more neutral. This effect can be observed in both photoluminescence and reflection measurements. This additional above-bandgap illumination is usually realized with the excitation laser for the photoluminescence study. During reflectivity measurements, apart from the incident light in the range corresponding to exciton structures, the sample should be illuminated additionally above the barrier energy gap – to increase the carrier density in the QW. Moreover, the carrier concentration can be estimated based on the ratio of the amplitudes of the spectral lines corresponding to the neutral X   and the charged X+   exciton present in the optical spectrum measured in a zero magnetic field. That allows one to simultaneously change and probe the carrier concentration in the nanostructure on an optical basis.
@@
II Sample and characterization
The studied structure was single 10 nm (Cd, Mn)Te QW grown by molecular beam epitaxy (MBE) between symmetrical 50-nm (Cd,Mg)Te barrier layers with Mg content of about 20 %. It was prepared on GaAs substrate with a thick buffer containing (3500 nm) CdTe layer with an additional 2000 nm (Cd,Mn)Te layer (Mn content the same as in the barriers). The Mg content in the barrier layer was determined by the reflectance measurements from the bandgap energy of the barrier Waag et al. (1993). The Mn2+ content in the QW material close to 0.3 % was chosen to ensure sufficiently large Zeeman splitting while preserving narrow excitonic lines. The exact value was confirmed by fitting the Brillouin function Gaj et al. (1994) to the exciton Zeeman shift obtained from reflectance measurements in a magnetic field, see Fig.1(b).
Generally, (Cd,Mn)Te/(Cd,Mg)Te QWs are naturally p-type, even without intentional doping. In such a case, hole gas originates from background doping of the barrier and/or from the surface states, which can be especially influential for the QW placed near the surface Maślana et al. (2003). Indeed, the presented samples contain hole gas in the QW. The sign of the carriers can be confirmed by using measurements of the photoluminescence signal in the magnetic field and observation of the spin singlet-triplet transition Kossacki et al. (2004); Łopion et al. (2020). The same group of samples was presented in Ref. Łopion et al., 2020, where their optical properties (along with the field of the spin singlet-triplet transition) were studied as a function of external illumination.
@@ II Sample and characterization
In the magnetic field, both exciton lines show a giant Zeeman shift. This shift is proportional to the magnetization of the Mn2+system, as probed by band carriers (Gaj and Kossut (2010); gajRelationMagnetoopticalProperties1993). It is well described by the modified Brillouin function, which exhibits a strong dependence on the temperature. Thus, one can consider how introducing carriers into the system affects the thermalization of magnetic ions excited by microwave radiation in the ODMR experiment. Figure 1(b) shows the measurement of giant Zeeman splitting as a function of the magnetic field with low carrier density (without additional illumination). The modified Brillouin function was fitted to the splitting of exciton lines X   in a magnetic field. From that fit, the temperature of the system of polarizing ions is estimated at 1.73 K. An analogous measurement was carried out for the increased carrier density – the effective temperature of the ions increased slightly in this situation, up to 1.81 K.
Figure 1(c) shows reflection spectra measured with the increased carrier density for different magnetic fields. In one circular polarization, the X+ line disappears. The positively charged exciton consists of two holes with opposite spins. The optical transition X+ associated with the photocreation of a hole with a specific spin depends on the prior existence of a hole with an opposite spin Huard et al. (2000). Hence, the oscillator strength of the charged exciton X+ is proportional to the concentration of holes in the opposite spin subband. By tracking the changes in the relative intensity of the reflection lines X   and X+, it can be seen that in this case, the hole gas is completely polarized already in the field of about 0.1 T, Fig. 1(d). The following function, derived from Boltzmann distribution of holes between two spin levels, was fitted to the intensity changes in the field presented in the graph:
@@
where $A_{\mathrm{max}}$ is the maximum intensity of a given line, $\Delta_{v,Z}$ giant Zeeman splitting of the valence band, $T_{\mathrm{h}}$ is the carriers’ temperature. This results in a temperature of the carriers slightly higher than the temperature of the ion system without microwave radiation. In the present paper, all the ODMR measurements were conducted in the magnetic fields above 0.1 T, thus in the fully polarized hole gas regime. In the next parts, we show that the carrier gas stays in this regime even after the absorption of the MW radiation during the ODMR experiments.
Figure 1: (a) Difference in the reflection spectrum in zero magnetic fields, in the absence and the presence of additional illumination with a blue LED. Two excitonic features are marked – neutral exciton (X) and positively charged exciton (X+). (b) Zeeman splitting of the neutral exciton line for the unilluminated situation (black points) with the fitted modified Brillouin function (gray curve) and the sample subjected to additional illumination (increased carrier density) - yellow points and yellow curve, respectively. The temperature of the ion system was a fitting parameter and was set to 1.73 K in the non-illuminated situation and 1.81 K in the illuminated situation. The black curve represents the ODMR signal obtained without the blue LED backlight for the frequency of microwave radiation $f$ = 12.9 GHz, the red curve symbolizes the modified Brillouin function for the remaining parameters exactly as in the case of the gray curve but the temperature set to 2.68 K. Inset: enlarged view of the region showing the detailed ODMR signal. (c) The dependence of the reflection spectrum for different magnetic fields with increased carrier density; (d) Relative intensity of the lines of the neutral X   and positively charged exciton X+ depending on the magnetic field with fitted relationships – according to the formula (1).
@@ III ODMR - setup and method of the measurements
The transitions between Mn2+ energy levels due to absorption of the applied microwave (MW) radiation lead to a decrease in magnetization. This can be evidenced by the decrease of the giant Zeeman shift. The effect can also be expressed in terms of increased effective spin temperature, which will be discussed later in this work. All the ODMR measurements presented in this work were conducted in the pulsed mode of microwave radiation and optical excitation to minimize non-resonant thermal drifts. Here, we define the ODMR signal as the difference in the energy position of the excitonic line in the reflectance spectrum for microwave and light pulses overlapping (ON) and shifted (OFF). The duration of the microwave pulse is about a few milliseconds. A representative ODMR signal as a function of magnetic field (ODMR spectrum) for fixed microwave frequency is present in Fig. 1(b). Careful study of ODMR spectra can provide information about strain, what was discussed in detail by Bogucki et al. (2022). Moreover, the ODMR spectrum can provide information about the impact of interactions with carriers, which is the core of the work presented in this paper.
The amplitude of the ODMR signal (the shift of the excitonic line) corresponds directly to the temperature of the magnetic ions, which creates net magnetization. After the sample is excited with pulsed microwave radiation of frequency of 12.9 GHz, a decrease in Zeeman splitting for a resonant magnetic field of 0.46 T can be observed, see Fig.1(b). This reduction in Zeeman splitting corresponds to an increase in the temperature of the ion system to 2.7 K as shown by the red curve in the figure.
An important detail is that the measurement conditions are set to avoid non-resonant signals. It can be evidenced by the fact that outside the resonance pattern, the position of the excitonic line with and without the microwaves is the same – see inset of Fig. 1(b). During the experiment, the overall temperature increase results from the resonant and non-resonant heating of the magnetic ions system. The measurements were conducted in the pulsed mode, in the timescale where we can assume that the non-resonant heating is constant. In Fig. 2(a), the ODMR spectra obtained with increasing microwave excitation power are presented. The signals reveal two distinct components: the resonant ODMR spectrum, which appears within a specific range of magnetic fields where resonance conditions are met, and a non-resonant signal that persists across the entire range of magnetic fields. This non-resonant offset is attributed to the non-resonant heating of the magnetic ion system via the crystal lattice or carrier gas. In Section V, we show that adding carriers to the sample (e.g., by additional illumination) impacts non-resonant heating in the presence of microwave radiation. Thus we can assume that most of the non-resonant heating is connected to the absorption of the MW in the carrier gas.
The excitation microwave power input can be directly connected to the magnetic ions’ temperature derived from the Zeeman splitting, as shown in Fig. 2(b). The input power was set to -10 dBm for further microwave radiation measurements to minimize the non-resonant heating. For such an excitation power, the ODMR signal is relatively good, and the temperature derived from the Zeeman splitting (1.81$\pm$0.01 K) is only slightly elevated in comparison to the temperature without microwaves (1.73$\pm$0.01 K), Fig. 1(b).
@@ IV Shape and position of the ODMR signal
The ODMR spectrum has a specific shape governed directly by the structure of Mn2+ energy levels. They can be described by the spin HamiltonianQazzaz et al. (1995); Bogucki et al. (2022); Abragam (1961):
@@
Figure 2: (a) ODMR signal power dependence, both amplitude of the signal and the offset are increasing with the applied microwave power (b) Temperature derived from Zeeman splitting change vs. microwave excitation power dependence. Full Giant Zeeman splitting for this sample, bath temperature, and the magnetic field are about 3 meV. That corresponds to the temperature marked with a red line – derived from Zeeman splitting measured without microwave radiation. The microwave excitation power in the ODMR measurements was set to -10 dBm to give a significant ODMR signal and avoid non-resonant signals, which are represented by an increase in the offset.
@@
Figure 3: (a) Reflection spectrum of the sample measured at temperature $\sim$1.6 K in magnetic field $B=0.463$ T with microwave radiation resonant for this field (12.9 GHz) at in the "ON" and "OFF" states, the change of position is visible on both visible exciton complexes; (b) The ODMR signal as the difference in the positions of the energy lines observed in the reflection spectrum in the presence and absence of microwave radiation at the frequency $f$ = 12.9 GHz depending on the value of the applied external magnetic field; the signal observed on the line of the charged exciton X+ is shifted ($\Delta B_{\mathrm{res}}$ = 14 mT) towards lower magnetic fields compared to the signal obtained on the neutral exciton X.
$H_{\mathrm{Mn^{2+}}}=g_{\mathrm{Mn}}\mu_{\mathrm{B}}\boldsymbol{\mathrm{B}}\cdot\widehat{\boldsymbol{\mathrm{S}}}+A\widehat{\boldsymbol{\mathrm{I}}}\cdot\widehat{\boldsymbol{\mathrm{S}}}+\\ +\frac{a}{6}\left(\widehat{S}_{x}^{4}+\widehat{S}_{y}^{4}+\widehat{S}_{z}^{4}-\frac{S(S+1)(3S^{2}+3S-1)}{5}\right)+\\ +D\left(\widehat{S}_{z}^{2}-\frac{S(S+1)}{3}\right),$ (2)
@@
where $\widehat{\boldsymbol{\mathrm{S}}}$ and $\widehat{\boldsymbol{\mathrm{I}}}$ are shorthand for $\widehat{\boldsymbol{\mathrm{S}}}\otimes\mathbb{1}$,$\mathbb{1}\otimes\widehat{\boldsymbol{\mathrm{I}}}$ respectively. The first term of the Hamiltonian describes the Zeeman splitting of the energy levels in a magnetic field due to the electron d-shell spin ($S$). $g_{\text{Mn}}$ is the g-factor of the manganese ion. The second term describes hyperfine splitting, where $\widehat{\boldsymbol{\mathrm{I}}}$ denotes the manganese ion nuclear spin operators. Mn2+  ion has electronic spin S = 5/2 and nuclear spin I = 5/2. The third term for the level-splitting part of the Hamiltonian is due to the placement of the ion in the crystal field with the symmetry Td.
@@
The last term refers to the zero-field splitting related to lowering the symmetry of the crystal and corresponds to the uniaxial stress associated with deformation in the $z$ direction. The strain-induced axial-symmetry parameter $D$ is a component characteristic for a given sample, resulting from the parameters of growth, composition, and thickness of individual layers. $D$ can be determined from the mismatch of the material lattice constants of the quantum well $a_{\mathrm{QW}}$ and the buffer $a_{\mathrm{buf}}$ by determining the strain $\varepsilon=\frac{a_{\mathrm{QW}}-a_{\mathrm{buf}}}{a_{\mathrm{QW}}}$. Finally, the $D$ parameter Qazzaz et al. (1995) can be determined: $D=-\frac{3}{2}G_{11}\left(1+\frac{2C_{12}}{C_{11}}\right)\varepsilon$, where $G_{11}$ = 72.272.272{.}272.2 neV and $\frac{2C_{12}}{C_{11}}=0.7$Bogucki et al. (2022); Greenough and Palmer (1973).
@@
The electronic and nuclear spin coupling for S and I equal to 5/2 results in 36 energetic states of Mn2+ ion. In a zero magnetic field, the splitting of those states corresponds mainly to the strain ($D$ parameter). If the magnetic field is non-zero but small (B<0.05T), the evolution of the states is complicated, and the eigenstates of the Hamiltonian are still not eigenstates of the $S_{z}$ operator. The Zeeman term dominates Hamiltonian for higher magnetic fields, and the energy levels begin to have a well-defined spin projection to the magnetic field axis ($z$ direction). Considering the selection rules for magnetic dipole transitions ($\Delta S_{z}=\pm 1$, $\Delta I_{z}=0$), the ODMR spectrum contains 6 sets with 5 transitions in each.
@@
The shape of the ODMR spectrum depends on the separation and the population of energy levels participating in microwave absorption. The intensity of a single transition corresponds to the occupation of the states between which the transition occurs—thus, it is strictly connected to the temperature of the magnetic ions. Qualitatively, the features observed at the higher magnetic field correspond to the transitions between lower energy levels than features observed at the lower magnetic field.
@@
In the case without additional illumination, the carrier density is low, and the neutral exciton transition dominates the reflectance spectrum. When additional illumination is set on the sample, the carrier density is increased. That is evidenced in the reflectance spectrum, which shows another excitonic transition in the energy below X: charged exciton (X+). Both excitonic lines observed in the reflectance are sensitive to microwave radiation, Fig. 3(a). We observe that both X and X+ lines move towards lower energies when the light and microwave pulses coincide in time ("ON" state) relative to the position of the lines when the light and microwave pulses are significantly temporarily separated ("OFF" state).
@@
Here, we focus on the observed differences in the ODMR signal obtained as a change in the position of neutral and charged exciton optical features. Generally, the neutral and charged excitonic complexes have different natures and respond differently to experimental variables. However, in the ODMR measurements presented in this work, we can treat both excitonic features as probes for the system of magnetic ions, which is the main subsystem absorbing microwave radiation. With this approach, there should be no difference if the magnetic ions are examined using neutral, charged, or any other excitonic line, as the probed system (magnetic ions) is the same. However, this is not the case in the shown studies. We can observe differences in the ODMR signal obtained as the position change of the neutral excitonic feature and the charged one. Figure 3(b) shows the ODMR signal measured as a shift of both X and X+ lines. In the ODMR signal measured as a change of the position of the neutral exciton line, many narrow features can be seen. Those are related to the transitions between different states of the manganese ion split by the magnetic field, hyperfine interaction, and strain. At the same time, we notice that such a structure is not visible in the signal monitored on the charged exciton line, and the shape of the signal seems smoother. Moreover, the maximum of the ODMR signal for X+ line is shifted relative to the maximum of the signal measured on the neutral exciton towards lower magnetic fields. These differences suggest that the two lines probe different populations of Mn2+  ions, which exhibit different paramagnetic resonance broadenings and shifts.
In the next parts, we show that the resonance shift observed on the charged state line results from interactions between magnetic ions and charge carriers. The $\Delta B_{\mathrm{res}}$ is connected to the density of carriers. Here, and in a similar situation in the following parts of the paper, the center of the resonance for the excitonic state is calculated as the center of mass of the ODMR spectrum (ODMR signal vs. magnetic field). That approach gives us precise information about the position of the resonance when it is more complicated in shape. However, it is important for the calculation to subtract the offset of the ODMR spectrum correctly. A similar result can be obtained by fitting a peak function (e.g., Gaussian) to the data.
@@
It should be noted here that both signals are acquired simultaneously, as they are extracted from the same optical spectrum. The ODMR signal obtained by observing the X line appears unaffected by the interaction with the charge carriers, while that obtained by observing the X+ line is apparently modified. This result suggests that the absorption associated with the neutral exciton occurs in different, spatially separated regions than the absorption associated with the charged exciton. At the same time, the spot of light used for reflectance measurements covers and probes both types of areas of different local carrier densities. The sample is relatively homogenous regarding the optical response in the scale above the laser spot size ($\sim$100 $\mu m$). That suggests the discussed inhomogeneities of the carrier density distribution are relatively small, below the spatial resolution of this optical measurement.